# One Model for All Tasks: Leveraging Efficient World Models in Multi-Task Planning

### v1

@@ A.4 Common Settings and Hyperparameters
Weight Decay 0.0001
Batch Size 256
Discount Factor ($\gamma$) 0.997
Training Context Length ($H$) 10
World Model
Num. Transformer Layers 4 (Atari-8, DMC, Jericho), 8 (Atari-26)
@@
Unroll Steps 5
Num. Experts (per MoE layer) 8 + 1 (shared)
@@
Inference Context Length ($H_{\text{infer}}$) 4
Replay Buffer
Replay Buffer Size 100,000
Loss Function Weights. The weights for the loss function are configured based on the action space type of the environment, as detailed in Table 8.
@@
Loss Term Discrete (Atari, Jericho) Continuous (DMC)
Observation Loss 10.0 10.0
@@
A.5 Training Framework and Computational Cost
All experiments were implemented using the LightZero [^4] https://github.com/opendilab/LightZero framework. All multi-task learning experiments were conducted on a single node equipped with 8x NVIDIA A100 (80G) GPUs. The approximate training times are shown in Table 9.
Table 9: Approximate training times for multi-task learning experiments.
Task Approximate Training Time
@@ B.1 Benchmark Setup
We use the Arcade Learning Environment (ALE) (Bellemare et al., 2013b) as our evaluation platform. Our experiments adopt two distinct training paradigms: multitask learning (MT), where a single model is trained to master all tasks simultaneously, and single-task learning (ST), where a dedicated model is trained independently for each task.
For our primary evaluation, we compare our multitask model, ScaleZero, with a robust single-task baseline using individually trained UniZero models. We use ScaleZero to denote the multitask (MT) setting and UniZero to denote the single-task (ST) setting. The MT/ST notation is sometimes omitted for brevity where the context is unambiguous. This comparison strategy is chosen because the multitask version of UniZero performs poorly, and there is currently no widely accepted standard baseline for multitask learning on the Atari benchmark. Consequently, UniZero ST serves as a more challenging and meaningful point of reference.
Our experiments are conducted in two stages: first, we perform model ablation studies on the Atari8 multitask benchmark (comprising Alien, Boxing, ChopperCommand, Hero, MsPacman, Pong, RoadRunner, and Seaquest); subsequently, we evaluate the final model’s performance on the full Atari26 multitask benchmark.
All game observations are preprocessed into 64x64 grayscale images, with 4-frame stacking used for state input. We employ standard environment wrappers, including sticky actions (p=0.25) and a frame skip of 4(Mnih et al., 2013). Performance is measured by the Human-Normalized Score (HNS), and we report both the mean and median scores. To ensure a fair comparison, all environmental settings are aligned with those in the UniZero paper (Pu et al., 2024).
B.2 Supplemental Plasticity Analysis
@@ B.2.1 Training Dynamics and Representational Collapse in UniZero
While Figure 1 in the main text shows performance degradation and worsening plasticity metrics for UniZero on complex tasks, Figure 7 illustrates that on four relatively simpler tasks from the Atari8 set, UniZero’s training process and corresponding metrics remain stable. This suggests a correlation between task difficulty and plasticity degradation.
To further probe the underlying cause of this degradation, we introduce the concept of effective rank (Dohare et al., 2024) to quantify the dimensionality of the model’s representation space. A low effective rank suggests representational collapse, where the model fails to maintain diverse features across tasks. Formally, consider a matrix $\bm{\Phi}\in\mathbb{R}^{n\times m}$ with singular values $\sigma_{k}$ for $k=1,2,\ldots,q$, where $q=\min(n,m)$. Let $p_{k}=\sigma_{k}/\sum_{i=1}^{q}\sigma_{i}$ be the normalized singular values, which form a probability distribution. The effective rank of $\bm{\Phi}$ is then defined as the exponential of the entropy of this distribution:
@@
The effective rank is a continuous measure ranging from 1 to the conventional rank of the matrix, providing a more nuanced view of the dimensionality of the space spanned by the matrix’s columns or rows. In the context of a neural network, the effective rank of a hidden layer’s activation matrix measures the effective number of dimensions in its representation space. A low effective rank indicates that the layer’s output can be generated by a small number of units, implying that many neurons are not contributing useful, independent information. In our experiments, we estimate the effective rank from a minibatch of 256 sequence samples before training on each task.
Figure 6 illustrates this concept, providing a diagnostic view of the representation’s dimensionality. The observation that a standard baseline model can exhibit a low effective rank provides quantitative evidence for the representational collapse hypothesis.
@@
Figure 6: Supplementary analysis of return and representation effective rank. This figure provides additional diagnostics for the baseline model shown in Figure 1, illustrating the relationship between game return (Left) and representation effective rank (Right). The performance degradation, observed as a drop in return, is largely accompanied by a corresponding decline in the representation effective rank. This provides further evidence that the performance collapse is linked to a degeneration of the model’s representation space.
@@
Figure 7: Training dynamics of UniZero on four other tasks from the Atari8 benchmark. This figure complements Figure 1 of the main text, showing the performance curves, dormant neuron ratio of the Transformer backbone, and latent state norm for Boxing, MsPacman, RoadRunner, and Alien.
B.2.2 Complete Metrics for Design Space Ablation Study
Figure 8 provides the complete data corresponding to Figure 3 in the main text, detailing the plasticity metrics for different model variants across all Atari8 tasks. Compared to the Baseline and different variants, the MoE architecture more effectively suppresses the growth of the dormant neuron ratio and the inflation of the latent state norm during training. This suggests that the MoE architecture is better at maintaining representational diversity and plasticity.
Figure 8: Complete plasticity metric analysis for the ScaleZero design space exploration. This figure details the evolution of performance (Return), Dormant Ratio, and Latent Norm for different model variants across all 8 tasks in the Atari8 benchmark.
Intrinsic Connections and Performance Implications
@@ B.2.2 Complete Metrics for Design Space Ablation Study
The observed correlations highlight a critical challenge in multi-task learning for standard Transformer architectures: representational collapse. The mechanism of this phenomenon is as follows: the model struggles to maintain distinct representations across tasks, causing its feature space to degenerate into a "one-size-fits-all" low-dimensional space, which is suboptimal for any individual task. This collapse is quantitatively indicated by a low effective rank, a problem diagnosed in baseline models (Section 1). The contraction of the representation space renders many neurons redundant, leading to a rising dormant neuron ratio. To compensate for this impoverished representational capacity, the model is forced to scale the magnitude of remaining active representations, which in turn causes the inflation of the latent norm. Ultimately, this collapsed and impoverished representation space lacks the capacity to model the diversity of multiple tasks, resulting in poor overall performance.
@@
In contrast, the MoE variant is designed to mitigate this very problem. By using its expert-based routing to preserve specialized representational subspaces, it addresses the symptoms of representational collapse. As shown in Figure 8, the MoE architecture effectively promotes neuron activity (suppressing the dormant ratio) and stabilizes representation magnitudes (controlling latent norm inflation), leading to superior multitask performance.
B.3 Performance on the Atari26 Multitask Benchmark
As shown in Table 1 of the main text, the mean HNS of the single multitask ScaleZero (MT) model on the Atari26 benchmark surpasses the average HNS of 26 independently trained UniZero single-task (ST) models. This result suggests that the multitask model achieves positive knowledge transfer across tasks. We hypothesize that this is due to the MoE architecture maintaining the network’s learning plasticity, enabling the transfer of general priors (e.g., object motion, collision detection) from simpler tasks to more complex ones.
@@ B.3 Performance on the Atari26 Multitask Benchmark
However, the multitask model’s median HNS is lower than the single-task baseline. This is primarily influenced by a few tasks with high exploration difficulty or unique mechanics (e.g., PrivateEye, Pitfall), indicating that negative transfer between heterogeneous tasks is not fully resolved in the current architecture. Figure 9 provides a comparison of the learning curves for all tasks.
@@
Figure 9: Performance comparison between ScaleZero (MT) and UniZero (ST) on the Atari26 benchmark. The figure shows the learning curves (mean and 95% confidence interval) for the multitask ScaleZero and the single-task UniZero baselines across 26 Atari games.
@@ C.1 Benchmark Setup
We use the DeepMind Control Suite (DMC) Tunyasuvunakool et al. (2020) as the evaluation benchmark for continuous control tasks. The experiments were conducted on a suite of 18 tasks (e.g., walker_walk, cheetah_run).
Following a similar methodology to our Atari experiments, we compare a single multitask ScaleZero (MT) model against a baseline of 18 single-task UniZero (ST) models, each trained independently on one task. All ScaleZero variants discussed in this section are multitask (MT) models.
The model input consists of low-dimensional state vectors provided by the environment, and the action space is a continuous vector normalized to the [-1, 1] range. The evaluation metric is the Average Return over 3 random seeds.
@@
As shown in Figure 10, the multitask ScaleZero model achieves returns comparable or superior to the single-task UniZero baseline on the majority of DMC tasks. This result demonstrates that the ScaleZero architecture is applicable to control problems involving high-dimensional continuous state and action spaces.
We hypothesize that this performance stems from the model learning shared physics priors and control primitives across different robotic morphologies. The expert modules within the MoE architecture may specialize in different types of dynamics (e.g., balancing, locomotion), which are then combined by the gating network based on the current task and state.
@@
Figure 10: Learning curves of ScaleZero (MT) vs. UniZero (ST) on the DMC18 benchmark. This figure compares the performance of the multitask ScaleZero and single-task UniZero models on 18 DMControl tasks. Solid lines represent the mean performance over 3 random seeds, and the shaded area indicates the 95% confidence interval.
C.3 Efficiency Evaluation of ScaleZero-DPS
Figure 11 compares the per-task learning curves of ScaleZero-DPS and the standard ScaleZero model to illustrate the difference in sample efficiency. The Figure shows that ScaleZero-DPS (orange line) exhibits a faster learning rate on most tasks, achieving equivalent performance levels with fewer environment steps. The asterisks (*) in the figure mark tasks that were identified as "solved" by the DPS policy and for which training was terminated early, showing the dynamic resource allocation mechanism in effect.
@@ C.3 Efficiency Evaluation of ScaleZero-DPS
Figure 11: Performance and efficiency comparison of ScaleZero-DPS vs. standard ScaleZero on DMC18. This figure shows the learning curves (return vs. environment steps) for both models across 18 tasks. An asterisk (*) indicates that the task’s training was terminated early by the DPS policy.

### v2

@@ A.5 Hyperparameter Settings
Weight Decay 1e-4
World Model (Transformer)
@@
Number of Heads 8
@@
Our ScaleZero model incorporates Mixture-of-Experts layers within its Transformer backbone for the Atari and DMC benchmarks to enhance model capacity and specialization. In contrast, the model for the Jericho benchmark does not use MoE layers. When MoE is used, each MoE layer consists of 8 specialist experts and 1 shared expert, with the router selecting the top-1 expert per token. The final loss is a weighted sum of several components, with weights varying between discrete and continuous action space environments, as detailed in Table 8.
@@
Loss Term Discrete Action Spaces (Atari, Jericho) Continuous Action Spaces (DMC)
@@
Observation Loss 10.0 10.0
A.5.1 Atari-Specific Settings
For Atari, observations are $64\times 64\times 3$ images processed by a ViT encoder. Training is conducted for 400,000 total environment steps.
Table 9: Key hyperparameters for Atari experiments.
Hyperparameter Value
General Training
Effective Global Batch Size 512
Discount Factor ($\gamma$) 0.997
World Model
Num. Transformer Layers 2 (Atari-8), 4 (Atari-26)
Training Context Length ($H$) 10
Encoder Type ViT
MCTS
Inference Context Length ($H_{\text{infer}}$) 4
Replay Buffer
Replay Buffer Size 500,000
Replay Ratio 0.25 (4-layer model), 0.5 (8-layer model)
A.5.2 DMC-Specific Settings
For the DMC suite, state vectors are processed by an MLP encoder. Training runs for 400,000 total environment steps.
Table 10: Key hyperparameters for DMC experiments.
Hyperparameter Value
General Training
Effective Global Batch Size 512
Discount Factor ($\gamma$) 0.99
Frame Skip 4 (8 for Pendulum)
World Model
Num. Transformer Layers 4
Training Context Length ($H$) 5
Encoder Type MLP
Num. Sampled Actions 20
MCTS
Inference Context Length ($H_{\text{infer}}$) 2
Replay Buffer
Replay Buffer Size 1,000,000
Replay Ratio 0.25
A.5.3 Jericho-Specific Settings
For the text-based Jericho suite, a pre-trained language model (BAAI/bge-base-en-v1.5) is used as the text encoder. Training runs for 500,000 total environment steps.
Table 11: Key hyperparameters for Jericho experiments.
Hyperparameter Value
General Training
Effective Global Batch Size 256
Discount Factor ($\gamma$) 0.997
World Model
Num. Transformer Layers 2
MoE Layers None
Training Context Length ($H$) 10
Encoder Type Text (BAAI/bge-base-en-v1.5)
MCTS
Inference Context Length ($H_{\text{infer}}$) 4
Replay Buffer
Replay Buffer Size 500,000
Replay Ratio 0.1
A.6 Training Framework and Computational Cost
All multitask learning experiments were conducted using one open-sourced framework on a single node equipped with 8x NVIDIA A100 (80G) GPUs. The approximate training times are shown in Table 12. We promise to open source the relevant code after the review process is completed.
A.6.1 Multi-Task Training Implementation Details
Our multi-GPU, multitask training framework is implemented based on PyTorch’s Distributed Data Parallel (DDP) (Li et al., 2020). The core logic is as follows:
  • Static Task Partitioning: The entire set of tasks (e.g., 26 Atari games) is statically partitioned among the available GPUs (ranks). Each GPU is assigned an exclusive subset of tasks to manage.
  • Per-Task Resource Allocation: On each GPU, dedicated resources, including a data collector, an evaluator, and a replay buffer, are instantiated for each assigned task.
  • Aggregated Batch Construction: During each training iteration, every GPU constructs a single training batch by sampling data segments from all replay buffers it manages. This creates a heterogeneous batch containing experiences from multiple tasks.
  • Synchronized Gradient Update: This aggregated batch is processed by the shared model to compute a single gradient. The DDP framework then averages the gradients across all GPUs, and the resulting synchronized gradient is used to perform a single optimization step on the shared model parameters.
This approach enables the joint optimization of a single model using data streams from multiple environments in a distributed and efficient manner.
Table 12: Approximate training times for multitask learning experiments in Section 4.
Benchmark Approximate Training Time
@@ B.1 Benchmark Setup
This section details the experimental setup for the Atari benchmarks discussed in Section 4.1. We use the Arcade Learning Environment (ALE) (Bellemare et al., 2013) as our evaluation platform.
Our evaluation protocol compares our multitask model, ScaleZero (MT), against a strong single-task baseline comprising individually trained UniZero (ST) models. This comparison is chosen because the multitask version of UniZero exhibits significant performance degradation, and a standard, widely-accepted baseline for online MTRL on the full Atari suite is not available. The UniZero ST baseline thus provides a more challenging and meaningful point of reference for evaluating positive knowledge transfer.
Our experiments proceed in two stages: first, we conduct model design and ablation studies on the Atari8 (multitask) benchmark (a subset of 8 games: Alien, Boxing, ChopperCommand, Hero, MsPacman, Pong, RoadRunner, and Seaquest). Second, we evaluate the final ScaleZero architecture on the full Atari26 (multitask) benchmark.
For all games, observations are preprocessed into 64x64 RGB images. The state input to the model is a single frame, resulting in an input tensor of shape 3x64x64 without any frame stacking. We use standard environment wrappers, including sticky actions (p=0.25) and a frame skip of 4 (Mnih et al., 2013). Performance is measured by the Human-Normalized Score (HNS), with both mean and median scores reported to assess overall capability. All environmental settings are strictly aligned with those in the UniZero paper (Pu et al., 2024) and DI-engine (Niu et al., 2021) to ensure a fair comparison.
B.2 Supplementary Analysis of Plasticity and Representational Collapse
This section provides a deeper empirical analysis of the plasticity loss and representational collapse phenomena diagnosed in Section 4.
@@ B.2.1 Training Dynamics and Representational Collapse in UniZero
While Figure 1 in the main text highlights performance degradation and plasticity loss for UniZero on complex tasks, Figure 7 shows that on four simpler tasks from the Atari8 set, UniZero’s training remains stable. This suggests a strong correlation between task difficulty and the onset of plasticity degradation.
To quantitatively diagnose the cause of this degradation, we measure the effective rank (Dohare et al., 2024) of the model’s representation space. A low effective rank indicates representational collapse, where the model fails to maintain diverse features required for multitask learning. For a matrix $\bm{\Phi}\in\mathbb{R}^{n\times m}$ with normalized singular values $p_{k}$, the effective rank is the exponential of the Shannon entropy of its singular value distribution:
@@
In our context, we compute the effective rank of a hidden layer’s activation matrix from a minibatch of 256 sequence samples. As shown in Figure 5, the performance collapse in the baseline model is directly accompanied by a decline in the representation’s effective rank, providing quantitative evidence for the representational collapse hypothesis.
@@
Figure 5: Analysis of representation effective rank. This figure supplements the diagnosis in Figure 1, showing the relationship between game return (Left) and representation effective rank (Right) for the baseline model. The sharp drop in performance correlates strongly with a decline in effective rank, substantiating the claim that performance failure is linked to a collapse in the dimensionality of the model’s representation space.
Figure 6: Correlation of performance with plasticity metrics for model variants. On two challenging tasks (ChopperCommand and Seaquest), the performance collapse of baseline models directly correlates with catastrophic plasticity loss, evidenced by a sharp rise in the dormant neuron ratio and latent norm. In stark contrast, the MoE-based model (ScaleZero) sustains high returns by maintaining a near-zero dormant ratio and a stable latent norm. This confirms that MoE’s superior performance stems directly from its success in mitigating the diagnosed plasticity issues. Full results for all 8 tasks are in Figure 8.
@@
Figure 7: Training dynamics of UniZero on four simpler Atari8 tasks. Complementing Figure 1 from the main text, this figure shows stable performance curves, dormant neuron ratios, and latent state norms for UniZero on Boxing, MsPacman, RoadRunner, and Alien, suggesting plasticity loss is task-difficulty dependent.
B.2.2 Mechanism of Representational Collapse
@@
The correlations observed in our experiments reveal a critical failure mode in standard multitask architectures. The model, struggling to reconcile conflicting task demands, forces its feature space to degenerate into a low-dimensional, "one-size-fits-all" subspace, which is suboptimal for any individual task. This is the mechanism of representational collapse, quantitatively diagnosed by a low effective rank. This contraction of the feature space renders many neurons redundant, leading to a rising dormant neuron ratio. To compensate for this impoverished representational capacity, the model is forced to scale the magnitude of the few remaining active features, causing an inflation of the latent norm. Ultimately, this collapsed representation lacks the capacity to model task diversity, resulting in poor performance.
@@
In contrast, ScaleZero’s MoE architecture is designed to mitigate this exact problem. By routing tasks to specialized expert sub-networks, it preserves distinct representational subspaces. As shown in Figure 8, the MoE architecture effectively maintains neuron activity (low dormant ratio) and stabilizes representation magnitudes (controlled latent norm), thereby preventing representational collapse and enabling superior multitask performance.
Figure 8: Complete plasticity metric analysis for the ScaleZero design space exploration. This figure provides the full data for the ablation study, detailing the evolution of performance (Return), Dormant Ratio, and Latent Norm for different model variants across all 8 tasks in the Atari8 multitask benchmark. The MoE-based model consistently outperforms others by maintaining superior plasticity metrics.
B.3 Full Performance on the Atari26 Benchmark
This section provides the complete learning curves corresponding to the summary results in 1(a). As reported in the main text, the single multitask ScaleZero (MT) agent achieves a higher mean HNS than the average of 26 specialized UniZero (ST) agents. This result demonstrates that ScaleZero achieves net positive knowledge transfer, which we attribute to the MoE architecture’s ability to maintain plasticity and transfer general priors (e.g., object physics) from simple to complex tasks.
@@ B.3 Full Performance on the Atari26 Benchmark
However, the median HNS of ScaleZero is lower than the ST baseline. This is primarily influenced by poor performance on a few hard-exploration or mechanically unique games (e.g., PrivateEye, Pitfall), indicating that negative interference is not fully eliminated. The complete learning curves for all 26 games are shown in Figure 9.
@@
Figure 9: Learning curves for ScaleZero (MT) vs. UniZero (ST) on the Atari26 benchmark. This figure shows the full per-game learning curves (mean and 95% confidence interval) for the multitask ScaleZero agent compared against 26 single-task UniZero baselines.
@@ C.1 Benchmark Setup
This section details the experimental setup for the DeepMind Control (DMC) Suite (Tunyasuvunakool et al., 2020), which serves as our benchmark for continuous control as presented in Section 4.1. Experiments were conducted on a suite of 18 tasks (e.g., walker_walk, cheetah_run), following the experimental setup of Pu et al. (2024). Following the methodology of our Atari experiments, we compare a single ScaleZero (MT) model against a baseline of 18 individually trained UniZero (ST) models. Model inputs are low-dimensional state vectors from the environment, and the action space is a continuous vector normalized to [1,1][-1,1]. Performance is measured as the average return over 8 evaluation episodes, with results averaged across 2 random seeds.
@@
As summarized in 1(b) and detailed in Figure 10, the multitask ScaleZero model achieves performance competitive with, and often superior to, the single-task UniZero baseline. Notably, ScaleZero achieves a higher median score, indicating robust generalist performance across the majority of tasks rather than excelling on only a few. This validates that the architectural principles of ScaleZero are effective for complex, state-based continuous control problems. We hypothesize this success stems from the MoE layers learning to specialize in shared physical priors and control primitives (e.g., balancing vs. locomotion), which are then composed by the router based on the task.
@@
Figure 10: Learning curves of ScaleZero (MT) vs. UniZero (ST) on the DMC18 benchmark. This figure provides the full per-task learning curves, comparing the performance of the single multitask ScaleZero agent against 18 specialized UniZero baselines. Solid lines represent the mean performance over 2 random seeds, and the shaded area indicates the 95% confidence interval.
C.3 Efficiency Evaluation of Dynamic Parameter Scaling (DPS)
This section provides the detailed learning curves that substantiate the efficiency gains of Dynamic Parameter Scaling (DPS), as reported in Section 4.2. Figure 11 compares the per-task sample efficiency of ScaleZero-DPS against the standard ScaleZero model.
The curves demonstrate that ScaleZero-DPS (orange) learns significantly faster on most tasks, achieving target performance levels with fewer environment interactions. This validates the claim that DPS leads to a substantial reduction in sample complexity. The asterisks (*) denote tasks where training was terminated early by the DPS policy upon reaching a "solved" threshold, visually confirming the dynamic allocation of computational resources away from mastered tasks.
@@ C.3 Efficiency Evaluation of Dynamic Parameter Scaling (DPS)
Figure 11: Sample efficiency comparison of ScaleZero-DPS vs. standard ScaleZero on DMC18. This figure shows the learning curves (return vs. environment steps) for both models across 18 tasks. The faster rise of the ScaleZero-DPS curves (orange) illustrates its superior sample efficiency. An asterisk (*) indicates that the DPS policy halted training for that task, demonstrating the dynamic resource allocation mechanism in action.
C.4 Mechanistic Analysis of Dynamic Parameter Scaling
To understand the mechanisms by which Dynamic Parameter Scaling (DPS) achieves its sample efficiency, we conduct a granular analysis of the agent’s internal parameter dynamics. This section presents two complementary visualizations: a time-series plot showing the continuous evolution of key metrics (Figure 12), and a suite of "importance matrices" providing discrete snapshots at critical training junctures (Figure 13).
C.4.1 Setup
Time-Series Dynamics (Figure 12): This figure provides a macroscopic view of the training process. The top panel correlates high-level metrics—normalized performance scores and the number of solved tasks—with the progression of training stages. The bottom panel offers a view into the internal mechanics by tracking the average learnable importance (‘alpha‘ scale) of LoRA adapters. It specifically compares the behavior of adapters in the first Transformer layer (Layer 0, solid lines) versus the last (Layer 3, dashed lines), highlighting the emergence of different dynamic behaviors at different model depths.
Stage-Wise Importance Matrices (Figure 13): This multi-panel figure offers a more granular, event-based perspective. Each matrix is an "importance matrix" where a cell at ‘(row r, column c)‘ shows the average ‘alpha‘ scale of the adapter introduced in Stage ‘c‘, as measured at the end of Stage ‘r‘’s training. This reveals how the model re-evaluates the contribution of all historical adapters after each phase of targeted learning. To provide a comprehensive view, this analysis is presented for:
  • The overall average across all layers and adapter types (Panel a).
  • A per-layer breakdown for Layer 0 and Layer 3 to show hierarchical differences (Panels b-c).
  • A per-type breakdown for adapters applied to the Query, Key, Value, and output Projection linear layers to reveal functional specialization (Panels d-g).
C.4.2 Results and Analysis
The analysis of these figures reveals several key dynamics of the DPS-managed learning process. The time-series plot (Figure 12) illustrates the staged training protocol. In this particular experiment, the stage transitions (vertical dashed lines) are triggered by a pre-allocated iteration budget, consistent with the method’s budget-based trigger mechanism. These capacity expansions occur as overall performance and the number of solved tasks increase over the course of training. The divergence in ‘alpha‘ scales between Layer 0 and Layer 3 suggests that the model applies different update strategies to shallow versus deep layers as training progresses.
The importance matrices in Figure 13 provide further detail on this behavior:
  • Overall Strategy (Panel a): The data is consistent with a dual mechanism of knowledge retention and plasticity. The importance scale of the foundational adapter from Stage 0 (column 0) is not static; it increases from 1.0 to 1.097 by the end of training. This suggests that the model increases the relative contribution of its initial knowledge as it encounters more complex tasks. Concurrently, the diagonal values show how the model focuses on the newest adapter during each stage, which is then re-weighted in subsequent stages as its knowledge is integrated.
  • Hierarchical Differentiation (Panels b vs. c): A notable difference is observed between Layer 0 and Layer 3. In Layer 0 (Panel b), the importance of the initial adapter (cell (4,0)) exhibits high stability with moderate growth (to 1.058), indicating its role as a stable foundation for low-level features. In contrast, Layer 3 (Panel c) shows more pronounced dynamic re-weighting. The foundational adapter’s importance grows more significantly (to 1.099), suggesting that higher-level, abstract representations rely heavily on and amplify this core knowledge. The greater variance in Layer 3’s alpha values indicates that later layers are more involved in adaptive, task-specific strategy modifications.
  • Functional Differentiation by Layer Type (Panels d-g): The analysis reveals distinct behaviors based on which linear layer within the attention mechanism an adapter modifies.
    • Query (Q) and Key (K) Adapters (Panels d, e): Adapters applied to the Q and K layers, which produce the representations for calculating attention scores, generally exhibit importance scales that increase over time. For instance, the Stage 0 Key adapter’s scale reaches 1.104. This suggests a continuous strengthening of the mechanisms that determine token-to-token relevance.
    • Value (V) and Projection (Proj) Adapters (Panels f, g): In contrast, adapters for the V layer (which provides the content to be aggregated) and the output Projection layer (which combines attention head outputs) show more complex dynamics, including suppression (alpha < 1.0). For example, the Value adapter from Stage 3 is down-weighted to 0.889 after Stage 4. This behavior may indicate a more cautious integration of new value-based information or a re-balancing of attention head outputs to integrate new skills without disrupting existing ones.
In summary, the DPS method facilitates a structured learning process. The model appears to preserve and amplify foundational knowledge, primarily in early layers, while enabling dynamic, adaptive strategy-switching in later layers. Furthermore, it modulates the importance of adapters based on their specific function within the self-attention mechanism (Query/Key vs. Value/Projection). This combination of knowledge retention and targeted, hierarchical plasticity offers a mechanistic explanation for the method’s observed sample efficiency.
Figure 12: Internal dynamics of ScaleZero-DPS during training. (Top) Performance and task-solving progress across stages. (Bottom) Average ‘alpha‘ scales for adapters in the first (Layer 0, solid) and last (Layer 3, dashed) layers, showing the emergence of hierarchical specialization.
(a) Overall Average
(b) Layer 0 Only
(c) Layer 3 Only
(d) Query Adapters
(e) Key Adapters
(f) Value Adapters
(g) Projection Adapters
Figure 13: Evolution of Adapter Importance Across Training Stages and Model Components. Each matrix displays the learned importance (average ‘alpha‘ scale) of adapters. A cell at ‘(row r, col c)‘ represents the importance of the adapter from Stage ‘c‘ after Stage ‘r‘’s training is complete. (a) The overall average shows a clear pattern of knowledge preservation (column 0) and targeted plasticity (diagonal). (b-c) The layer-specific breakdown reveals that foundational knowledge is more rigidly preserved in early layers (Layer 0) while later layers (Layer 3) are more dynamic and adaptive. (d-g) The type-specific breakdown shows functional differentiation in how adapters are utilized over time, reflecting their distinct roles in the attention mechanism.